# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v3

@@ IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
IV-F2 Evaluation on Deepfake Datasets
As we mention before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [47]. We report the results in Table IX. Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
TABLE XI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
@@ IV-F2 Evaluation on Deepfake Datasets
TABLE XII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
$N$ ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
@@
IV-G Ablation Study
@@
We conduct an ablation study in Table XI to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
@@
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in LSUN(ProGAN) dataset. With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper. As shown in Table XII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
@@ V Conclusion
In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks. For the task of AIGI detection, we propose a novel frequency-based post-train Bayesian attack, which explores the vulnerable regions in the frequency domain in a Bayesian manner. More broadly, our proposed post-train Bayesian attack strategy extends the Bayesian attack family. Extensive experiments have been conducted across models, generators and defense methods under both white-box and black-box settings. The following conclusions can be drawn from our investigation.
  • The current SOTA AIGI detectors are vulnerable to adversarial attacks, while the widely-used defense mechanism adversarial training cannot be directly applied to the AIGI detectors, which brings great challenges to build reliable and robust detectors.
  • AIGI detectors rely on gradient masking to mitigate the transferability of gradient-based attacks, but the effect of gradient masking can be easily circumvented by a simple random initialization strategy.
  • The attack success rate is positively associated with the detection accuracy. Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
These findings indicate that building an adversarial robust AIGI detector is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to this domain.
Acknowledgments
This work was supported in part by the NSF China (No. 62302139) and FRFCU-HFUT (JZ2023HGTA0202, JZ2023HGQA0101).

### v4

@@ IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
IV-G2 Evaluation on Deepfake Datasets
As we mention before, there is a big difference between AIGI detection and Deepfake detection. But we empirically found that FPBA is also effective for face forgery detection. To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [1]. We report the results in Table XIII. Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance. For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
TABLE XV: Attack success rate (%) under different frequency transforms on the ProGAN dataset.
Surrogate Transform CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Avg
CNNSpot FFT 98.75 98.15 76.05 95.70 19.60 47.40 50.85 94.45 72.62
Wavelet 99.20 98.10 71.20 94.55 19.30 46.90 50.60 92.65 71.56
DCT 98.90 98.00 76.40 95.90 19.80 48.00 51.50 94.80 72.91
MobileNet FFT 31.65 40.00 51.00 99.50 22.35 36.50 12.45 44.80 42.28
Wavelet 23.04 31.15 32.04 91.40 21.85 36.45 11.20 36.55 35.46
DCT 32.20 40.40 51.70 99.60 22.60 36.40 12.30 44.60 42.48
TABLE XVI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
@@ IV-G2 Evaluation on Deepfake Datasets
TABLE XVII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
NN ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
@@
IV-H Ablation Study
The Selection of Frequency Transforms
We conduct an ablation study comparing the performance of different frequency transforms, including Fast Fourier Transform (FFT), Wavelet Transform (Wavelet) and Discrete Cosine Transform (DCT). The results in Table XV shows that FFT, Walevet and DCT all show high attack performance, with DWT performing the best. So we use DCT by default.
@@ IV-H Ablation Study
We conduct an ablation study in Table XVI to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
@@
In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models. The adversarial examples were crafted from ResNet-50 in Synthetic LSUN(ProGAN) dataset. With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper. As shown in Table XVII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate. Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation. Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
@@ V Conclusion
In this paper, we investigate the adversarial robustness of AIGI detectors and propose a novel frequency-based post-train Bayesian attack that extends the Bayesian attack family. Through extensive experiments across models, generators, and defenses under both white-box and black-box settings, we draw the following conclusions: (1) State-of-the-art AIGI detectors remain highly vulnerable to adversarial attacks. Although adversarial training is effective in other domains, it cannot be directly applied in AIGI detection, posing major challenges for building robust detectors. (2) Many detectors rely on gradient masking to hinder the transferability of gradient-based attacks, yet the defense is easily bypassed with simple random initialization. (3) Attack success correlates positively with detection accuracy, suggesting that robustness evaluation on low-accuracy subsets is uninformative. Overall, our findings highlight that constructing adversarially robust AIGI detectors remains an open problem. We hope our work serves as both a benchmark and a call to further research in this critical area.